# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

### v2

@@ 4. Proposed Dependency Update Metrics
In this section, we first describe our design choices that fill the gaps presented in Section 3.2 for dependency update metrics measurement. Then, we provide an example of the nominal case using our metrics. We then provide a detailed methodology for defining the dependency update metrics in Section 4.3.
4.1. High-Level Overview
@@
We design our metrics to only consider direct dependencies since the package only has control over which version to use for the direct dependencies. We begin by examining each of the $<$package, dependency$>$ relationships from a temporal perspective. We then split each $<$package, dependency$>$ relations into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release). Our decision to split into intervals is to capture the benefits of floating version constraints without incentivizing this practice (Gap 1). [^1] To the best of our knowledge, no academic work quantifies optimal dependency specification (e.g., pinning vs floating version constraints) for balancing security benefits and the cost of maintenance. Therefore, our design does not discriminate between ways to specify dependency constraints (he_pinning_2025, ). Then, at each interval, the dependency constraint set by the package for the dependency is resolved with only the versions available at the beginning of the interval. We do this to make sure that the dependency resolution accounts for the historical version releases of the dependency.
@@
After splitting into intervals, we mark “updated” as ‘false’ if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise. Similarly, we mark “remediated” as ‘false’ if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3). We then aggregate the “updated” and “remediated” information of each $<$package, dependency$>$ to compute the MTTUdep and MTTRdep of each package (Gap 2).
4.2. Nominal Case With Our Metrics
@@
Table 2. Running example of codemod-cli package with one of it’s depndency simple-git.
row pkg pkg version dep dep constraint dep version dep highest rel. Interval start Interval end updated remediated
1 codemod-cli $0.4.0$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ false true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
82 codemod-cli $0.8.6$ simple-git ${}^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ false true
83 codemod-cli $0.8.7$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ false true
84 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ false true
85 codemod-cli $0.9.0$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ false false
86 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ false false
87 codemod-cli $0.9.1$ simple-git ${}^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ false false
88 codemod-cli $0.9.2$ simple-git ${}^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ true true
@@
In this section, we explain how to calculate the dependency update metrics. As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
@@
We split the dependency relations into multiple intervals based on the release of a new version of the package “codemod-cli” or dependency “simple-git”. In each interval, for resolving the dependency constraint by the package codemod-cli, we only consider the dependency versions available at the beginning of the interval (“interval start”). From $V0.4.0$ to $V0.8.6$ (row 1-82), codemod-cli has the dependency constraint simple-git ${}^{\wedge}1.130.0$ which resolves into $V1.132.0$. simple-git has the highest release $V2.13.1$ at $\operatorname{2020-07-16}$ (row 1) and $V3.2.6$ at $\operatorname{2022-03-03}$. As a result, codemod-cli does not have the highest available release of simple-git in these intervals (rows 1-82). In $V0.8.7$, codemod-cli updated the constraint for simple-git to ${}^{\wedge}2.48.0$. Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3 and rather stayed at major version 2. codemod-cli have ${}^{\wedge}2.48.0$ constraint for simple-git from $V0.8.7$ to $V0.9.1$. The false value in the “updated” column means that the package has an outdated version of the dependency in this interval.
@@
At the beginning of row 85, the resolved version of simple-git $V2.48.0$ was found vulnerable to four vulnerabilities (CVE-2022-24066 (CVE-2022-24066, ), CVE-2022-24433 (CVE-2022-24433, ), CVE-2022-25912 (CVE-2022-25912, ), and CVE-2022-25860 (CVE-2022-25860, )). In this example, we only consider CVE-2022-24433 as a vulnerability, which was fixed in $V3.3.0$, released on $\operatorname{2022-03-11}$. For this reason, we mark the intervals from rows 85-87 as “remediated” = false since codemod-cli has a vulnerable version of simple-git even though a fixed version is available. The false value in the “remediated” column means that the package has an outdated and vulnerable version of the dependency in this interval. In $V0.9.2$, codemod-cli bumped the constraint of simple-git to ${}^{\wedge}3.4.0$ which resulted in $V3.4.0$, a fixed version of the above vulnerability. So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked as true in the “remediated” column.
When computing the Time-To-Update (TTUdep) for dependency relation between codemod-cli and simple-git, we sum up the intervals marked false in the “updated” column. For codemod-cli, the TTUdep is the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87) which is $614$ days. Similarly, for computing Time-To-Remediate (TTRdep), we sum up the intervals marked false in the “remediated” column. For codemod-cli, the TTRdep is the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87) which is $11$ days.
@@ 4.3. Metrics Definitions
Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$).
MTTUdep of a package is the average aggregated time the package uses an outdated direct dependency version in its lifetime.
We have described how to calculate the TTUdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(2) ${TTU}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ updated=False}$
MTTUdep of package $p_{i}$ with $n$ direct dependencies is:
@@
Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$).
MTTRdep of a package is the average aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
We have described how to calculate the TTRdep for a $<$package, dependency$>$ relationship in Section 4.2. Formally speaking, TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($<p_{i},p_{j}>$ relationship) is:
(4) ${TTR}_{dep}(p_{i},p_{j})=\sum\operatorname{intervals\ with\ remediated=False}$
MTTRdep of package $p_{i}$ with $n$ direct historically vulnerable dependencies is:
@@
5. Methodology
In this section, we first present vulnerability, package metadata (versions and dependency relations), and package characteristics collection process. We then present a statistical testing process to verify if MTTUdep can be a proxy for MTTRdep (RQ1). We then present correlation tests and regression analysis to understand how package characteristics impact MTTUdep and MTTRdep (RQ1).
@@ 5.1.1. Vulnerability Information
We collect the CVE data for our chosen three ecosystems from osv.dev (osv-dev, ) for npm, PyPI, and Cargo packages on 2024-09-12. We rely on OSV since OSV aggregates CVE data from multiple sources (e.g., GitHub Security Advisories, PyPA, GoVulDB) (osv-data-sources, ) in one place. After downloading JSON formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability is fixed.
5.1.2. Package Metadata
@@
We collect the package-version data and dependency information for npm, PyPI, and Cargo packages from deps.dev on 2024-08-20, similar to other previous studies (lin_untrustide_2024, ; zahan_software_2023, ). In our dataset, we have initially $2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo packages. We collected data from deps.dev since it provides all the package versions and dependency information for our three chosen ecosystems. After data collection, we split each $<$package, dependency$>$ relation into multiple intervals. Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints with our added requirement of only using the available versions of the dependency before the interval start time.
5.1.3. Package Characteristics
@@
We examined several characteristics (from Saini et al. (saini_investigating_2020, )) for packages: the number of contributors; the number of dependencies and dependents; the number of version releases; the SourceRank score, and the number of forks and stars. On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these factors correlate with package’s MTTUdep and MTTRdep values. Libraries.io is used by other research as a data source (gu_self-admitted_2023, ; cao_towards_2023, ; he_migrationadvisor_2021, ). When counting the number of dependencies, we only use the number of dependencies of packages in their latest version. To ensure construct validity in downloading these factors, we manually inspected a sample of packages to verify the factors, and we found the downloaded factors to be accurate. Additionally, we computed the number of major versions and package ages for our packages and added them to our analysis.
5.2. Package Inclusion Exclusion Criteria
@@
We begin with an initial dataset of $3,000,103$ ($2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo) packages collected from deps.dev. Our first step is to apply two inclusion criteria: (1) the package must be at least two years old (operationalized by the difference between the first and last version release); and (2) the package must have at least one residual activity (e.g., one version release) in the last two years. Miller et al. (miller_understanding_2025, ) used two years of residual activity followed by two years of no maintenance as criteria to find out the abandoned packages. Our criteria are inspired by Miller et al. since we want to include the packages that are maintained. Moreover, “two years” is a commonly used standard adopted by other research to measure whether a package is maintained or not (li_comparison_2023, ). However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )) even if they are not abandoned. We then used our exclusion criteria: a package without any dependency should be excluded. Since our metrics characterize packages’ dependency update practice, packages without any dependencies do not fit into our study. The number of packages after this exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages). Out of these $163,207$ packages, $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages have at least one vulnerable dependency in their lifetime, supporting our initial motivation of lack of vulnerability data.

### v3

@@ 4. RQRQ1:: Novel Dependency Update Metrics
In this section, we first describe our design choices that fill the gaps presented in Section 3.2 and lead to our novel metrics design (MTTUdep and MTTRdep) for RQRQ1:: . Then, we provide an example case study using our metrics. We then provide a detailed methodology for defining our novel dependency update metrics in Section 4.3.
4.1. Intuition
@@
We design our metrics to only consider direct dependencies since the package only has control over which version to use for the direct dependencies. We begin by examining each of the <<package, dependency>> relationships from a temporal perspective. We then split each <<package, dependency>> relation into multiple intervals based on when the package or the dependency has a newer release (major, minor, or patch release). Our decision to split into intervals is to capture the benefits of floating version constraints without incentivizing this practice (Gap 1). [^2] To the best of our knowledge, no academic work quantifies optimal dependency specification (e.g., pinning vs floating version constraints) for balancing security benefits and the cost of maintenance. Therefore, our design does not discriminate between ways to specify dependency constraints (he_pinning_2025, ). Then, at each interval, the dependency constraint set by the package for the dependency is resolved with only the versions available at the beginning of the interval. We do this to make sure that the dependency resolution accounts for the historical version releases of the dependency.
@@
After splitting into intervals, we mark as “updated”=false if the resolved dependency version for that interval does not match the highest available version of the dependency at that time and ‘true’ otherwise. Similarly, we mark as “remediated”=false if the resolved dependency version is vulnerable with a fixed version being available at the beginning of that interval and ‘true’ otherwise (Gap 3). We then aggregate the “updated” and “remediated” information of each <<package, dependency>> to compute the MTTUdep and MTTRdep of each package (Gap 2). This aggregation involves factoring in how old each interval is and weighting based on that. With our weighting mechanism, recent intervals get near full weight, and older intervals’ weight drops exponentially to zero (Gap 4).
4.2. An Example With Our Metrics
@@
Table 2. Running example of codemod-cli package with one of its dependency simple-git.
row
pkg
pkg
version
dep
dep
constraint
dep
version
dep
highest rel.
Interval start Interval end
Age Of Interval
updated remediated
1 codemod-cli $0.4.0$ simple-git $^{\wedge}1.130.0$ $1.132.0$ $2.13.1$ $\operatorname{2020-07-16}$ $\operatorname{2020-07-17}$ 1495 false true
@@
82 codemod-cli $0.8.6$ simple-git $^{\wedge}1.130.0$ $1.132.0$ $3.2.6$ $\operatorname{2022-03-03}$ $\operatorname{2022-03-05}$ 899 false true
83 codemod-cli $0.8.7$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-05}$ $\operatorname{2022-03-08}$ 896 false true
84 codemod-cli $0.9.0$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.2.6$ $\operatorname{2022-03-08}$ $\operatorname{2022-03-11}$ 893 false true
85 codemod-cli $0.9.0$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-11}$ $\operatorname{2022-03-17}$ 887 false false
86 codemod-cli $0.9.1$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.3.0$ $\operatorname{2022-03-17}$ $\operatorname{2022-03-18}$ 886 false false
87 codemod-cli $0.9.1$ simple-git $^{\wedge}2.48.0$ $2.48.0$ $3.4.0$ $\operatorname{2022-03-18}$ $\operatorname{2022-03-22}$ 882 false false
88 codemod-cli $0.9.2$ simple-git $^{\wedge}3.4.0$ $3.4.0$ $3.4.0$ $\operatorname{2022-03-22}$ $\operatorname{2022-03-29}$ 875 true true
$\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$ $\ldots$
@@
In this section, we explain how to calculate the dependency update metrics, Time-To-Update (TTUdep) and Time-To-Remediate (TTRdep), for a <<package, dependency>> relationship. As a running example, we consider the dependency relation between codemod-cli and simple-git, as shown in Table 2.
@@
We split the dependency relations into multiple intervals based on the release of a new version of the package “codemod-cli” or dependency “simple-git”. In each interval, for resolving the dependency constraint by the package codemod-cli, we only consider the dependency versions available at the beginning of the interval (“interval start”). From $V0.4.0$ to $V0.8.6$ (row 1-82), codemod-cli has the dependency constraint simple-git $^{\wedge}1.130.0$ which resolves into $V1.132.0$. simple-git has the highest release $V2.13.1$ at $\operatorname{2020-07-16}$ (row 1) and $V3.2.6$ at $\operatorname{2022-03-03}$. As a result, codemod-cli does not have the highest available release of simple-git in these intervals (rows 1-82). In $V0.8.7$, codemod-cli updated the constraint for simple-git to $^{\wedge}2.48.0$. Even with this update, codemod-cli has not changed the constraint of simple-git to the available highest major version 3, and rather stayed at major version 2. codemod-cli have $^{\wedge}2.48.0$ constraint for simple-git from $V0.8.7$ to $V0.9.1$. “updated”=false indicates that the package has an outdated version of the dependency in this interval.
@@
At the beginning of row 85, the resolved version of simple-git $V2.48.0$ was found vulnerable to four vulnerabilities (CVE-2022-24066 (CVE-2022-24066, ), CVE-2022-24433 (CVE-2022-24433, ), CVE-2022-25912 (CVE-2022-25912, ), and CVE-2022-25860 (CVE-2022-25860, )). In this example, we only consider CVE-2022-24433 as a vulnerability, which was fixed in $V3.3.0$, released on $\operatorname{2022-03-11}$. For this reason, we mark the intervals from rows 85-87 as “remediated” = false since codemod-cli has a vulnerable version of simple-git even though a fixed version is available. “remediated”=false means that the package has an outdated and vulnerable version of the dependency in this interval. In $V0.9.2$, codemod-cli bumped the constraint of simple-git to $^{\wedge}3.4.0$, which resulted in $V3.4.0$, a fixed version of the above vulnerability. So codemod-cli has the highest available version of simple-git in this interval (row 88) and so marked “remediated”=true.
When computing the Time-To-Update (TTUdep) for the dependency relation between codemod-cli and simple-git, we sum up the intervals with “updated”=false with exponential weighting. With the weighting factor, TTUdep for codemod-cli becomes 2.4 days, which is less than naively computing the delta between $\operatorname{2020-07-16}$ to $\operatorname{2022-03-22}$ (rows 1 - 87), which is $614$ days. Since this period of outdated dependency (rows 1 - 87) occurred in the year 2022, the weighting factor ensures giving less emphasis on that. Similarly, TTRdep for codemod-cli (rows with “remediated”=false) becomes 3.67 days. Because of the weighting factor, our computed TTRdep =3.67 is lower than naively summing up the delta between $\operatorname{2022-03-11}$ to $\operatorname{2022-03-22}$ (rows 85 - 87), which is $11$ days. Since this period of intervals with “remediated”=false is older, these intervals are weighted accordingly in TTRdep. We formally define MTTUdep and MTTRdep in Section 4.3 and explain our design choices.
@@ 4.3. Metrics Definitions
Metric 1 ($\operatorname{\mathbf{Mean-Time-To-Update}}_{dep}$ : $\mathbf{MTTU}_{dep}$)
MTTUdep of a package is the weighted aggregated time the package uses an outdated direct dependency version in its lifetime.
We have described how to calculate the TTUdep for a <<package, dependency>> relationship in Section 4.2. Formally speaking, TTUdep of a package $p_{i}$ with considering only dependency $p_{j}$ ($p_{i},p_{j}>$ relationship) is defined in Equation 2.
(2) ${TTU}_{dep}(p_{i},p_{j})=\frac{\sum_{t}w_{t}d_{t}}{\sum_{t}w_{t}}$
MTTUdep for package $p_{i}$ with $n$ direct dependencies is defined in Equation 3.
@@
Here, $d_{t}$ indicates one interval duration with ‘updated’=false for $p_{i},p_{j}>$ relationship which ends at timestamp $t$. Also, $w_{t}=\exp(-\lambda a_{t})$ is the weight assigned for interval $d_{t}$ and $a_{t}$ is the age of the interval $d_{t}$. $\lambda$ is the decaying factor in this weighting function, and we set $\lambda=\frac{\ln(2)}{\tau}$. In this equation, $\tau$ is the half-life, and we set $\tau=2\ years$ since 2 years is used in literature to assess recent ongoing maintenance (miller_understanding_2025, ; li_comparison_2023, ). With this decaying weight, recent intervals approach full weight, while the weight of the older intervals decays exponentially to near zero. Our use of weighted average is inspired by similar other research (guzman_emotional_awareness_2013, ; ajiono_time_series_forecasting_2023, ; lee2016comparison, ; sikand_green_ai_2023, ).
We considered linear ($w_{t}=max(a)-a_{t}+\epsilon$), exponential ($w_{t}=\exp(-\lambda a_{t})$), and inverse ($w_{t}=\frac{1}{a_{t}+\epsilon}$) as the choices for the weighting function based on the criteria described by Ulan et al. (ulan_weighted_2021, ) for weighted quality scoring for software metrics. We opted for the exponential weighting function since exponential weighting is more responsive to recent data than linear weighting, and is more configurable and robust than inverse weighting (using $\tau$). For example, if a development team wants to consider 3 years as an appropriate half-life for their specific case, they can configure the weighted version by changing $\tau=3\ years$.
Metric 2 ($\operatorname{\mathbf{Mean-Time-To-Remediate}}_{dep}$ : $\mathbf{MTTR}_{dep}$)
MTTRdep of a package is the weighted aggregated time a package uses an outdated and vulnerable direct dependency version in its lifetime.
TTRdep of a package $p_{i}$ with considering only dependency $p_{j}$ is defined in Equation 4.
(4) ${TTR}_{dep}(p_{i},p_{j})=\frac{\sum_{t}w_{t}d_{t}}{\sum_{t}w_{t}}$
MTTRdep for $n$ direct dependencies is defined in Equation 5.
@@
Here, $d_{t}$ indicates one interval duration with ‘remediated’=false for $p_{i},p_{j}>$ relationship which ends at timestamp $t$. Also, $w_{t}=\exp(-\lambda a_{t})$ is the weight assigned for interval $d_{t}$ and $a_{t}$ is the age of the interval $d_{t}$.
Takeaway 1: Our design of MTTUdep and MTTRdep overcomes the limitations of existing dependency update metrics.
5. Empirical Study Methodology
In this section, we first present vulnerability, package metadata (versions and dependency relations), and the packages’ characteristics collection process to apply the metric implementation in the three ecosystems (RQRQ2:: ). After that, we present a statistical testing process to verify if MTTU can be a proxy for MTTR (RQRQ3:: ). Lastly, we present correlation tests and regression analysis to understand how package characteristics impact MTTU and MTTR (RQRQ4:: ).
@@ 5.1. Data Collection
Vulnerability Information. We collect the CVE data for our chosen three ecosystems from osv.dev (osv-dev, ) for npm, PyPI, and Cargo packages on 2024-09-12. We rely on OSV since OSV aggregates CVE data from multiple sources (e.g., GitHub Security Advisories, PyPA, GoVulDB) (osv-data-sources, ) in one place. After downloading JSON-formatted CVE data from OSV, we convert it into an SQL table that includes ecosystem, package name, CVE ID, version where the vulnerability was introduced, and version where the vulnerability was fixed.
@@
Package Metadata. We collect the package-version data and dependency information for npm, PyPI, and Cargo packages from deps.dev on 2024-08-20, similar to other previous studies (lin_untrustide_2024, ; zahan_software_2023, ). We chose the three ecosystems to have a diverse set of ecosystems, where npm is the largest, PyPI is the oldest, and Cargo is the newest among the major software ecosystems. In our dataset, we have initially $2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo packages. We collected data from deps.dev since it provides all package versions and dependency information for our three chosen ecosystems. After data collection, we split each <<package, dependency>> relation into multiple intervals. Since the dependency resolution can be complex and differs across ecosystems, we use deps.dev to perform the dependency resolution allowed by the dependency constraints, with our added requirement (only using the available versions of the dependency before the interval start time).
@@
Package Characteristics. We examined several characteristics (from Saini et al. (saini_investigating_2020, )) for packages: the number of contributors; the number of dependencies and dependents; the number of version releases; the SourceRank score, and the number of forks and stars. On 2025-01-11, we downloaded these characteristics of each package from libraries.io to understand if these characteristics influence a package’s MTTUdep and MTTRdep values. Libraries.io is used by other research as a data source (gu_self-admitted_2023, ; cao_towards_2023, ; he_migrationadvisor_2021, ). When counting the number of dependencies, we only use the number of dependencies of packages in their latest version. To ensure construct validity in downloading these characteristics, we manually inspected a sample of packages to verify the characteristics, and we found the downloaded characteristics to be accurate. Additionally, we computed the number of major versions and package ages for our packages and added them to our analysis.
5.2. Package Inclusion and Exclusion Criteria
@@
We begin with an initial dataset of $3,000,103$ ($2,603,314$ npm, $274,720$ PyPI, and $122,069$ Cargo) packages collected from deps.dev. Our first step is to apply two inclusion criteria: (1) the package must be at least two years old (operationalized by the difference between the first and last version release); and (2) the package must have at least one residual activity (e.g., one version release) in the last two years. Miller et al. (miller_understanding_2025, ) used two years of residual activity followed by two years of no maintenance as criteria to find out the abandoned packages. Our criteria are inspired by Miller et al., since we want to include the packages that are maintained. Moreover, “two years” is a commonly used standard adopted by other research to measure whether a package is maintained or not (li_comparison_2023, ). However, our package selection criteria might miss packages that are less than two years old or have had no activity in the last two years (e.g., feature complete packages (coelho2017modern, )), even if they are not abandoned. We then used our exclusion criteria: a package without any dependencies should be excluded. Since our metrics characterize packages’ dependency update practice, packages without any dependencies do not fit into our study. The number of packages after these exclusion criteria is $163,207$ ($117,129$ npm, $42,777$ PyPI, and $3,301$ Cargo packages), which will be the final set we use in our analyses. Out of these $163,207$ packages, $22,513$ ($17,263$ npm, $5,158$ PyPI, and $92$ Cargo) packages have at least one vulnerable dependency in their lifetime, supporting our initial motivation of lack of vulnerability data.